What is the equation of the line that passes through the point (3, 2) and has a slope of 4
Answer by nerdybill(7384)

What is the equation of the line that passes through the point (3, 2) and has a slope of 4
.
Plug given information into the "point-slope" form of a line:
y - y1 = m(x - x1)
y - 2 = 4(x - 3)
y - 2 = 4x - 12
y = 4x - 10 (this is what they're looking for in "slope-intercept" form)